# Oxygen Isotope Ratios in Hydrogen-Deficient Carbon Stars: A Correlation with Effective Temperature and Implications for White Dwarf Merger Outcomes

### v1

@@ 3.5 Sources of systematic errors
We now discuss the limitations of our analysis that are likely to introduce some systematic effects in the derived elemental abundances. We note that as these limitations have similar effects on the strengths of 12C16O and 12C18O lines, they should not significantly alter the oxygen isotope ratios which are the main focus of our work.
@@ 3.5.1 Model Grid
The sparse grid of H-deficient stellar models available in the literature is a significant limitation in any elemental abundance study of HdC stars. It is clear that the range of input elemental abundances available in the model-grid are not suitable for fitting the spectra of all HdC stars in our sample. For several stars, the derived abundances differ substantially from the input values assumed in the MARCS models. We cannot quantify the inaccuracies introduced by these deviations. A new set of HdC atmospheric models that span a wider range of elemental abundances is required to address this issue. The abundances derived in this paper will help inform the input abundance choices for new models.
The MARCS models used in our analysis were made with a fixed value of 9.5 that appears to be too high for the cooler RCB stars in our sample. We find that ranges from 8.7 to 10.3. Similarly, is observed to range from 7.2 in the RCB NSV11154 to 9.0 in the dLHdC C38, and the mean value of 8.1 indicates that the fixed input model of 8.8 is too large for most stars. The nitrogen abundances observed are well represented by the grid of models available in , and this work highlights the importance of generating a similar grid of models for and . These discrepancies from the input model abundances are likely to introduce errors in the absolute abundances we report.
As noted earlier, most of the stars in our sample show [Fe] . It is clear that the assumed solar metallicity in the MARCS models is not appropriate for most of these stars, and an expansion of the model grid to include sub-solar metallicities is necessary.
@@ 3.5.2 Dust Dilution
RCB stars show an infrared (IR) excess due to the presence of warm circumstellar dust (Feast et al., 1997; García-Hernández et al., 2009; Tisserand, 2012). This is expected to contribute an additional continuum flux component that has a veiling effect over photospheric absorption lines, resulting in the lines appearing weaker than they actually are. This effect is evident in the spectrum of the IRAS 1813+ shown in Figures 2 and 2. As discussed in Section 2, this star was in a decline of mag from its maximum luminosity, indicating that a fresh dust ejection episode from this star has severely diluted the photospheric spectrum. This phenomenon is also expected to weaken the lines of the other RCB stars in our sample by a lesser extent, but this is an effect that we are unable to correct for. It is likely that the derived absolute abundances for these stars are slightly underestimated due to this effect, but the oxygen isotope ratios should be unaffected.
@@ 3.5.4 Carbon Problem
One major issue encountered while studying HdC stars is the so-called ‘carbon problem’ that was extensively discussed by Asplund et al. (2000) in their detailed analyses of high-resolution optical spectra of warm (T > 6250 K) RCB stars. The issue arises because the photoionization of neutral carbon atoms, that are in a highly excited states in such warm atmospheres, is the dominant source of continuous opacity. This leads to a discrepancy between the predicted and the observed C I lines strength. The carbon abundances derived from C I absorption lines are systematically lower by a median value of 0.6 dex than the input abundances in the stellar atmospheric models (Asplund et al., 2000).
@@
To get around the ‘carbon problem’, Hema et al. (2012, 2017) used the C2 lines to measure the carbon abundance from their optical high resolution spectra. This technique appeared to be an excellent alternative. However the estimatees with C2 lines resulted in values systematically lower than the ones found with C I lines, implying that the ratio C/He in HdC stars’ atmosphere should be much lower that the ratio of 0.6% measured in EHe stars (Asplund et al., 2000; Pandey et al., 2006). Consequently, they noted that, if one considers that both groups of stars should have similar modes of formation, a version of the carbon problem with the C2 lines could also affect their measurements. Overall, using warm and cold HdC stars, they estimated that the carbon abundance should be lower for all HdC stars, and extrapolated low C/He ratios ranging between 0.03% and 0.3%. Furthermore, it is worth noting that, with their colder HdC stars sample, their measure of the carbon abundance from the C I lines and from the C2 lines agreed (their Table 6), indicating possibly a variation of the ‘carbon problem’ with effective temperature.
@@
Right: Abundances of 18O with N (right) for the dLHdC (blue) and RCB (orange) stars in our sample. We find that is linearly correlated with (black dashed line), suggesting that a fixed fraction () of the nitrogen is converted to 18O across all stars via -capture.
@@
García-Hernández et al. (2010) underlined that, depending on the HdC star effective temperature, the main contributor to the continuous opacity may change to free-free absorption from neutral helium when observing in the infrared K band, thus the ‘carbon problem’ could disappear at such wavelength. García-Hernández et al. (2009, 2010) measured the carbon abundance for 5 cold RCB, 2 cold dLHdC and 2 warm dLHdC (T > 6250 K) stars from high resolution infrared spectra. They reported not seeing the ‘carbon problem’ with any of the HdC stars using the C I lines, and the measured carbon abundances were consistent with the input model (C/He=1%), within a range of -0.3 to +0.4 dex. However, they observed a ’carbon problem’ with the cold HdC stars using the C2 lines. Their measured carbon abundance was about 0.8 dex higher than the input abundance.
@@
Similar to García-Hernández et al. (2009, 2010), we observed cold HdC stars in the infrared. We used the new C2 ExoMol linelist (Yurchenko et al., 2018) to estimate the carbon abundances from C2 lines. For RCB stars, our derived C abundances range between 8.7 and 9.2, which is 0.6 dex lower than the input abundance of 9.52 (C/He=1%). However, conversely, for dLHdCs, our measured C abundances exceed the input abundance by 0.4 to 0.8 dex, which is similar to the result reported by García-Hernández et al. (2009). For the two dLHdC stars we have in common, HD 137613 and HD 182040, we derived a C2 based carbon abundance that is about 0.5 dex lower than the one reported by García-Hernández et al. (2009). That is possibly explained by the use of different C2 line lists.
@@
An expanded set of models with a grid of input carbon abundances is required to quantify the extent of the carbon problem with C2 lines. As the main focus of our study is oxygen isotope ratios, we leave this analysis for a future study and report the carbon abundances derived with our current models in Table 4.
@@ 4 Discussion
The main goal of this paper was to measure precise oxygen isotope ratios for a large sample of RCB and dLHdC stars. Using our high-resolution spectra, we find that all dLHdC stars in our sample have oxygen isotope ratios lower by an order of magnitude than all RCB stars in our sample. All dLHdC stars in our sample have 16O/18O , while all RCB stars have 16O/18O . The coolest star in our sample RCB WISE J1942+ has a ratio as high as 92. Our results robustly confirm that dLHdC stars tend to have lower 16O/18O ratios than RCB stars – a trend suggested by previous analyses of medium resolution spectra (Karambelkar et al., 2022) and small samples of high-resolution spectra (García-Hernández et al., 2010). As noted previously, this difference could potentially be explained by requiring that dLHdCs and RCBs originate in white-dwarf binaries with distinct properties such as total masses or mass-ratios (see Tisserand et al. 2022; Karambelkar et al. 2022; Crawford et al. 2023 for details).
@@
Additionally, we note a correlation between the oxygen isotope ratios and the effective temperatures of HdC stars (Figure 5 (left)), with warmer stars having lower 16O/18O ratios than the cooler ones. The two coolest stars ( K) have oxygen-isotope ratios that are an order of magnitude larger than the warmer ( K) stars. This is the first observed trend of oxygen isotope ratios in RCB and dLHdC stars. This trend of oxygen isotope ratios with the effective temperatures of HdC stars has also been predicted by the recent Crawford et al. (2024) theoretical models. They find that the lower (), higher Mtot mergers generally produce colder stars with higher oxygen isotope ratios, while high , low Mtot mergers produce hotter stars with low oxygen isotope ratios. Fig. 5 (left) also shows the oxygen isotope ratios as a function of effective temperatures from the Crawford et al. (2024) model grid, highlighting a dependence on . It is promising that the models predict the same general trend seen in the observations. However, as noted in Crawford et al. (2024), the exact values of the oxygen isotope ratios are not reproduced by the models. We note that the models overpredict the observed oxygen isotope ratios of RCBs by a factor of and dLHdCs by two orders of magnitude. Crawford et al. (2024) suggest that the model oxygen isotope ratios can be reduced by adopting lower helium-shell burning temperatures (TSOF), but the lowest ratio achieved for dLHdC models was , which is still significantly larger than our observations. Additionally the amount of hydrogen in the He-WD or the extent of convection can also impact the oxygen isotope ratios (Munson et al., 2021; Zhang et al., 2014). Further modeling to reproduce the observed values of the isotope ratios will shed light on the role of the WD progenitor properties (in addition to Mtot and ) in determining the observed properties of RCB and dLHdC stars.
@@
We also note a positive correlation between the abundances of 14N and 18O in the HdC stars (Figure 5, right). This is consistent with the -capture reaction chain 14N18FO believed to be responsible for the formation of 18O during partial He burning (Clayton et al., 2005). A linear fit to the data in Figure 5 (right) gives a slope of and a y-intercept of dex, suggesting that the amount of 18O produced is proportional to the amount of 14N present, and that of 14N is converted to 18O across all HdC stars.
@@ 5 Conclusion
In this paper, we have analyzed high-resolution (R) K-band (2.25 – 2.48 m) spectra of six RCB and six dLHdC stars to determine their chemical abundances and oxygen isotope ratios. We fit the observed spectra to synthetic spectra constructed using a grid of MARCS models for HdC stars and precisely measured their 16O/18O ratios. We find that all six RCB stars in our sample have 16O/18O and all six dLHdC stars have 16O/18O , confirming previous suggestions that dLHdC stars have lower oxygen isotope ratios than RCB stars. Additionally, for the first time, we find a possible correlation between the oxygen isotope ratios and the effective temperatures of HdC stars, with colder stars having larger 16O/18O values than their hotter counterparts. This trend is consistent with recent simulations of white-dwarf mergers that explore the effects of total mass and mass ratios of WD binaries on their merger remnants. However, these models overpredict the 16O/18O values by two orders of magnitude, pointing to additional factors that could play an important role in setting the oxygen isotope ratios.
@@ 6 Acknowledgements
AM and PT thank Sergey Yurchenko from the ExoMol team for providing diatomic molecular line lists, and Bertrand Plez for the discussions we had during the course of this article. PT thanks the MARCS team in Uppsala (Sweden) and especially Kjell Eriksson for kindly providing a grid of hydrogen-deficient stellar models. AM thanks Nick Storm and Jeffrey Gerber for suggestions that led to the method we developed to fit isotope ratios. MMK and VB thank the Department of Science and Technology, Government of India, and Indian National Academy of Engineering for an inaugural Vaishvik Bharatiya Vaigyanik (VAIBHAV) fellowship.

### v2

@@ 3.5 Sources of systematic errors
We now discuss the limitations of our analysis that are likely to introduce some systematic effects in the derived elemental abundances. We note that as these limitations have similar effects on the strengths of 12C16O and 12C18O lines, they should not significantly alter the oxygen isotope ratios which are the main focus of our work (Asplund et al., 2000; García-Hernández et al., 2009).
@@ 3.5.1 Model Grid
The sparse grid of H-deficient stellar models available in the literature is a significant limitation in any elemental abundance study of HdC stars. It is clear that the range of input elemental abundances available in the model-grid are not suitable for fitting the spectra of all HdC stars in our sample. For several stars, the derived abundances differ substantially from the input values assumed in the MARCS models. A new set of HdC atmospheric models that span a wider range of elemental abundances is required to address this issue. The abundances derived in this paper will help inform the input abundance choices for new models.
The MARCS models used in our analysis were made with a fixed value of 9.5 that appears to be too high for the cooler RCB stars in our sample. We find that ranges from 8.7 to 10.3. Similarly, is observed to range from 7.2 in the RCB NSV11154 to 9.0 in the dLHdC C38, and the mean value of 8.1 indicates that the fixed input model of 8.8 is too large for most stars. The nitrogen abundances observed are well represented by the grid of models available in , and this work highlights the importance of generating a similar grid of models for and . As noted earlier, most of the stars in our sample show [Fe] . It is clear that the assumed solar metallicity in the MARCS models is not appropriate for most of these stars, and an expansion of the model grid to include sub-solar metallicities is necessary.
These discrepancies from the input model abundances are likely to introduce errors in the absolute abundances we report. While we cannot directly quantify these errors due to a lack of models, Asplund et al. (2000) tested the effects for R CrB. They generated a new model matching the values derived from a model that deviated from the abundances obtained by as much as 1.3 dex – comparable to the differences we observe. They found that the abundances derived from the new model were within 0.05 dex of the original values, suggesting that the derived abundances are robust to deviations in the input abundances.
@@ 3.5.2 Dust Dilution
RCB stars show an infrared (IR) excess due to the presence of warm circumstellar dust (Feast et al., 1997; García-Hernández et al., 2009; Tisserand, 2012). This is expected to contribute an additional continuum flux component that has a veiling effect over photospheric absorption lines, resulting in the lines appearing weaker than they actually are. This effect is evident in the spectrum of IRAS 1813+ shown in Figures 2 and 2. As discussed in Section 2, this star was in a decline of mag from its maximum luminosity, indicating that a fresh dust ejection episode from this star has led to a diluted photospheric spectrum, compounding the effects of the existing circumstellar dust shell. This dust shell is also expected to weaken the lines of the other RCB stars in our sample, but it is an effect that is difficult to estimate and correct for. We would require spectral energy distributions (SEDs) from the same epoch as the spectra, along with accurate models of the properties of the dust shell, which is beyond the scope of this work.
Instead, we attempt to quantify this effect for a few different stars using a simplified model. We examine the effect of a dust shell continuum flux that is added to the photospheric flux and estimate the true photospheric spectrum. We define the following quantities:
: Total observed flux.
, : Total fluxes contributed by the photosphere and dust shell, respectively.
, : Continuum flux levels of the photosphere and dust shell, respectively. In the absence of any dust lines in the wavelength range of interest, we set
: Total continuum flux level.
: Fractional contribution of the dust shell to the total observed continuum flux (i.e. ).
: Continuum-normalized observed and photospheric fluxes, respectively. These can be written as ;
Our continuum normalized fluxes measure and we wish to determine .
Writing the total observed flux as :
Rearranging this gives :
Finally, dividing by C to convert to a continuum-normalized flux and simplifying gives the final relation:
Given a value of , we can now determine the true, undiluted continuum-normalized photospheric flux. It is important to note that , like the individual continuum levels, is a wavelength-dependent quantity that we assume to be fixed for simplicity. We chose as a reasonable value to test, as we estimate this quantity to be in the 10 – 30 range for these stars in the K-band. We detect no trend of the recovered abundances with wavelength, indicating that the gradient in is much smaller than the line-to-line scatter, and that fixing is a reasonable approximation. After applying this correction we observed an increase in depth of all lines, with stronger lines being affected the most. For most of the species that we analyze, the lines appear to be far from saturation, and we see a change of less than dex for all the elemental abundances other than that of oxygen. In the case of oxygen, the 12C16O lines are approaching saturation, and the abundance of 16O is therefore affected strongly, with the correction causing an increase of up to 0.4 dex depending on the star. However, since the 12C18O lines are rather weak, the change in abundance of 18O is similarly small like the other elements. This leads to an increase of 0.45 – 0.55 dex in which is dominated by the 16O fraction increasing, and an accompanying increase in the 16O/18O ratio by roughly a factor of 2. This indicates that the circumstellar dust shell has a significant effect on estimates of the oxygen isotope ratio, and is a factor that must be studied carefully in future works that use near-IR spectra of RCB stars.
@@ 3.5.4 Carbon Problem
One major issue encountered while studying HdC stars is the so-called ‘carbon problem’ that was extensively discussed by Asplund et al. (2000) in their detailed analyses of high-resolution optical spectra of warm (T > 6250 K) RCB stars. The issue arises because the photoionization of neutral carbon atoms, that are in highly excited states in such warm atmospheres, is the dominant source of continuum opacity. This leads to a discrepancy between the predicted and the observed C I lines strength. The carbon abundances derived from C I absorption lines are systematically lower by a median value of 0.6 dex than the input abundances in the stellar atmospheric models (Asplund et al., 2000).
@@
To get around the ‘carbon problem’, Hema et al. (2012, 2017) used the C2 lines to measure the carbon abundance from their optical high resolution spectra. This technique appeared to be an excellent alternative. However the estimates with C2 lines resulted in values systematically lower than the ones found with C I lines, implying that the ratio C/He in HdC stars’ atmosphere should be much lower that the ratio of 0.6% measured in Extreme Helium (EHe) stars (Asplund et al., 2000; Pandey et al., 2006). Consequently, they noted that, if one considers that both groups of stars should have similar modes of formation, a version of the carbon problem with the C2 lines could also affect their measurements. Overall, using warm and cold HdC stars, they estimated that the carbon abundance should be lower for all HdC stars, and extrapolated low C/He ratios ranging between 0.03% and 0.3%. Furthermore, it is worth noting that, with their colder HdC stars sample, their measure of the carbon abundance from the C I lines and from the C2 lines agreed (their Table 6), indicating possibly a variation of the ‘carbon problem’ with effective temperature.
@@
Right: Abundances of 18O with N (right) relative to Fe for the dLHdC (blue) and RCB (orange) stars in our sample. We find that () - A(Fe)) is linearly correlated with (black dashed line), suggesting that a fixed fraction () of the nitrogen is converted to 18O across all stars via -capture.
@@
García-Hernández et al. (2010) underlined that, depending on the HdC star effective temperature, the main contributor to the continuum opacity may change to free-free absorption from neutral helium when observing in the infrared K band, thus the ‘carbon problem’ could disappear at such wavelengths. García-Hernández et al. (2009, 2010) measured the carbon abundance for 5 cold RCB, 2 cold dLHdC and 2 warm dLHdC (T > 6250 K) stars from high resolution infrared spectra. They reported not seeing the ‘carbon problem’ with any of the HdC stars using the C I lines, and the measured carbon abundances were consistent with the input model (C/He=1%), within a range of -0.3 to +0.4 dex. However, they observed a ‘carbon problem’ with the cold HdC stars using the C2 lines. Their measured carbon abundance was about 0.8 dex higher than the input abundance.
@@
Similar to García-Hernández et al. (2009, 2010), we observed cold HdC stars in the infrared. We used the new C2 ExoMol linelist (Yurchenko et al., 2018) to estimate the carbon abundances from C2 lines. For RCB stars, our derived C abundances range between 8.7 and 9.2, which is 0.6 dex lower than the input abundance of 9.52 (C/He=1%). As noted in Section 3.5.2, correcting for dilution by circumstellar dust could increase by about 0.1 dex, which would reduce this gap slightly. However, conversely, for dLHdCs, our measured C abundances exceed the input abundance by 0.4 to 0.8 dex, which is similar to the result reported by García-Hernández et al. (2009). For the two dLHdC stars we have in common, HD 137613 and HD 182040, we derived a C2 based carbon abundance that are about 0.5 dex lower than the ones reported by García-Hernández et al. (2009). That is possibly explained by the use of different C2 line lists.
@@
Lastly, we are also able to estimate the carbon abundance from C I lines (A(C)* in Table 4) that are visible in the spectra of all the dLHdC stars in our sample but missing from the RCBs (likely due to a combination of lower carbon abundances and their cooler temperatures favoring C2 formation). We found these to be systematically lower than the values obtained using C2 lines, with the difference ranging from -0.42 to -0.13 with an average value of -0.23. We note that the uncertainties on these differences are up to 0.2 – 0.3 dex, so all the stars appear to be consistent with each other. This appears to disagree with the results reported by Hema et al. (2012), who observed the C2 abundances to be lower than C I for warm RCB stars using optical spectra. This discrepancy could again be related to a variation of the ‘carbon problem’ with effective temperature and/or wavelength.
An expanded set of models with a grid of input carbon abundances is required to quantify the extent of the carbon problem with C2 lines, and verify the hypothesis that the differences in the C/He ratios of RCB and dLHdC stars could be related to their formation scenarios. As the main focus of our study is oxygen isotope ratios, we leave this analysis for a future study and report the carbon abundances derived with our current models in Table 4.
@@ 4 Discussion
The main goal of this paper was to measure precise oxygen isotope ratios for a large sample of RCB and dLHdC stars. Using our high-resolution spectra, we find that all dLHdC stars in our sample have oxygen isotope ratios lower by an order of magnitude than all RCB stars in our sample. All dLHdC stars in our sample have 16O/18O , while all RCB stars have 16O/18O . The coolest star in our sample RCB WISE J1942+ has a ratio as high as 92. As discussed in Section 3.5.2, 16O/18O estimates may be underestimated by a factor of 2 for RCBs if the circumstellar dust shell contributes to 20% of the observed K-band flux. This would further widen the 16O/18O gap between RCB and dLHdC stars. Our results robustly confirm that dLHdC stars tend to have lower 16O/18O ratios than RCB stars – a trend suggested by previous analyses of medium resolution spectra (Karambelkar et al., 2022) and small samples of high-resolution spectra (García-Hernández et al., 2010). As noted previously, this difference could potentially be explained by requiring that dLHdCs and RCBs originate in white-dwarf binaries with distinct properties such as total masses or mass-ratios (see Tisserand et al. 2022; Karambelkar et al. 2022; Crawford et al. 2023 for details).
@@
Additionally, we note a correlation between the oxygen isotope ratios and the effective temperatures of HdC stars (Figure 5 left), with warmer stars having lower 16O/18O ratios than the cooler ones. The two coolest stars ( K) have oxygen-isotope ratios that are an order of magnitude larger than the warmer ( K) stars. This is the first observed trend of oxygen isotope ratios in RCB and dLHdC stars. This trend of oxygen isotope ratios with the effective temperatures of HdC stars has also been predicted by the recent Crawford et al. (2024) theoretical models. They find that the lower (), higher Mtot mergers generally produce colder stars with higher oxygen isotope ratios, while high , low Mtot mergers produce hotter stars with low oxygen isotope ratios. Fig. 5 (left) also shows the oxygen isotope ratios as a function of effective temperature from the Crawford et al. (2024) model grid, highlighting a dependence on . It is promising that the models predict the same general trend seen in the observations. However, as noted in Crawford et al. (2024), the exact values of the oxygen isotope ratios are not reproduced by the models. We note that the models overpredict the observed oxygen isotope ratios of RCBs by a factor of (this could potentially be bridged by the effects of dilution, as noted in Section 3.5.2) and dLHdCs by two orders of magnitude. Crawford et al. (2024) suggest that the model oxygen isotope ratios can be reduced by adopting lower helium-shell burning temperatures (TSOF), but the lowest ratio achieved for dLHdC models was , which is still significantly larger than our observations. Additionally the amount of hydrogen in the He-WD or the extent of convection can also impact the oxygen isotope ratios (Munson et al., 2021; Zhang et al., 2014). Further modeling to reproduce the observed values of the isotope ratios will shed light on the role of the WD progenitor properties (in addition to Mtot and ) in determining the observed properties of RCB and dLHdC stars.
@@
We also note a positive correlation between the abundances of 14N and 18O in the HdC stars (Figure 5, right). This is consistent with the -capture reaction chain 14N18FO believed to be responsible for the formation of 18O during partial He burning (Clayton et al., 2005). A linear fit to the data in Figure 5 (right) gives a slope of and a y-intercept of dex, suggesting that the amount of 18O produced is proportional to the amount of 14N present, and that of 14N is converted to 18O across all HdC stars. Dust shell dilution does not play a significant role here as the increase in these abundances was found to be dex, which is further mitigated by our use of abundance ratios.
@@ 5 Conclusion
In this paper, we have analyzed high-resolution (R) K-band (2.25 – 2.48 m) spectra of six RCB and six dLHdC stars to determine their chemical abundances and oxygen isotope ratios. We fit the observed spectra to synthetic spectra constructed using a grid of MARCS models for HdC stars and precisely measured their 16O/18O ratios. We find that all six RCB stars in our sample have 16O/18O and all six dLHdC stars have 16O/18O , confirming previous suggestions that dLHdC stars have lower oxygen isotope ratios than RCB stars. In our investigation of the possible effects of dilution caused by circumstellar dust in RCBs, we showed that the oxygen isotope ratios of RCB stars could be underpredicted by a factor of 2 if the dust emission contributes to 20% of the observed K-band flux, highlighting the need for careful characterization of this dust in future studies of RCBs that rely on infrared spectra. Additionally, for the first time, we find a possible correlation between the oxygen isotope ratios and the effective temperatures of HdC stars, with colder stars having larger 16O/18O values than their hotter counterparts. This trend is consistent with recent simulations of white-dwarf mergers that explore the effects of total mass and mass ratios of WD binaries on their merger remnants. However, these models overpredict the 16O/18O values by two orders of magnitude, pointing to additional factors that could play an important role in setting the oxygen isotope ratios.
@@ 6 Acknowledgements
AM and PT thank Sergey Yurchenko from the ExoMol team for providing diatomic molecular line lists, and Bertrand Plez for the discussions we had during the course of this article. PT thanks the MARCS team in Uppsala (Sweden) and especially Kjell Eriksson for kindly providing a grid of hydrogen-deficient stellar models. AM thanks Nick Storm and Jeffrey Gerber for suggestions that led to the method we developed to fit isotope ratios. MMK and VB thank the Department of Science and Technology, Government of India, and Indian National Academy of Engineering for an inaugural Vaishvik Bharatiya Vaigyanik (VAIBHAV) fellowship. This work has made use of the VALD database, operated at Uppsala University, the Institute of Astronomy RAS in Moscow, and the University of Vienna. The observations presented in this paper were obtained under the Visiting Astronomer Program at the Infrared Telescope Facility, which is operated by the University of Hawaii under contract 80HQTR24DA010 with the National Aeronautics and Space Administration. This publication makes use of data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), which is a joint project of the Jet Propulsion Laboratory/California Institute of Technology and the University of California, Los Angeles. NEOWISE is funded by the National Aeronautics and Space Administration. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain.
7 Software and Data availability
The high-resolution spectra of all HdC stars and the software used for abundance and isotope ratio measurements in this paper are available at this github repository : https://github.com/advaitmehla/HdC-high-res.